OpenStudy (anonymous):

Could someone please help me with sqrt(3) tan2x=0

OpenStudy (lgbasallote):

what do you want to do with it?

OpenStudy (anonymous):

solve for x

OpenStudy (lgbasallote):

\[\sqrt 3 \tan (2x) = 0\] first step divide both sides by \(\sqrt{3}\) what do you get?

OpenStudy (y2o2):

x = 0 or x = 180

OpenStudy (anonymous):

so tan2x=-1/sqrt(3)

OpenStudy (lgbasallote):

nope. \[\tan (2x) =\frac {0}{\sqrt3}\] now simplify this

OpenStudy (anonymous):

oh dear! sorry i didnt realise! The question is sqrt(3) tan2x+1=0

OpenStudy (anonymous):

sorry!

OpenStudy (lgbasallote):

haha..then you're right \[\tan(2x) =-\frac{1}{\sqrt 3}\]

OpenStudy (lgbasallote):

now second step. take the arctangent of both sides

OpenStudy (anonymous):

ok so pi/6?

OpenStudy (lgbasallote):

remember that negative sign...and it's not yet over

OpenStudy (anonymous):

how does the negative affect it?

OpenStudy (lgbasallote):

well it doesnt affect that much

OpenStudy (anonymous):

how do you mean?

OpenStudy (lgbasallote):



OpenStudy (anonymous):

oh ok so when you say take the arctangent, where do you go from there?

OpenStudy (lgbasallote):

so \[2x = -\frac{\pi}{6}\] now you divide both sides by 2

OpenStudy (anonymous):

ok so x= =pi/12

OpenStudy (anonymous):

-pi/12